# DeepQuark: deep-neural-network approach to multiquark bound states

### v1

@@ Appendix B Details of the optimization process
The deep neural network (DNN) in the DeepQuark wave function consists of the input features $\boldsymbol{x}=\left(\boldsymbol{r}_{i},\left|\boldsymbol{r}_{i}-\boldsymbol{r}_{j}\right|,\alpha_{c},\alpha_{s},\alpha_{t}\right)$, four hidden fully connected layers and a one-dimensional output $f_{NN}(\boldsymbol{x})$. Here $\boldsymbol{r}_{i},\alpha_{c},\alpha_{s},\alpha_{t}$ are the spatial three-dimensional coordinates in the center of mass frame and the encoded vectors of color, spin and isospin coupled bases, respectively. The number of nodes in each hidden layer and the total number of variational parameters in the DNN for different systems are given in Table 5. We use larger DNNs for larger systems to ensure the flexibility of the wave function.
@@
Table 5: The number of nodes in each hidden layer and the total number of variational parameters in the DNN for different systems.
@@
In each iteration step during the optimization process, we use the Metropolis-Hastings Monte Carlo algorithm [54, 55] to generate sample points that are distributed according to the probability
[…]
@@
$P=\frac{\left|\psi_{\boldsymbol{\theta}}\left(R^{\prime},\alpha^{\prime}\right)\right|^{2}}{\left|\psi_{\boldsymbol{\theta}}(R,\alpha)\right|^{2}}.$ (13)
The energy expectation value $E_{\boldsymbol{\theta}}$ and its gradient with respect to the parameters $\nabla_{\boldsymbol{\theta}}E_{\boldsymbol{\theta}}$ can be estimated using these sample points.
[…]
@@
$E_{L}(R,\alpha)=\sum_{\alpha^{\prime}}\psi_{\boldsymbol{\theta}}(R,\alpha)^{-1}H_{\alpha\alpha^{\prime}}(R)\psi_{\boldsymbol{\theta}}(R,\alpha^{\prime}).$ (15)
@@
$\nabla_{\boldsymbol{\theta}}E_{\boldsymbol{\theta}}=2$ $\left[\frac{1}{N}\left(\sum_{n}E_{L}(R_{n},\alpha_{n})\nabla_{\boldsymbol{\theta}}\log\psi_{\boldsymbol{\theta}}(R_{n},\alpha_{n})\right)\right.-$ (16)
@@
$\boldsymbol{\theta}^{i+1}=\boldsymbol{\theta}^{i}-\eta(S+\epsilon I)^{-1}\nabla_{\boldsymbol{\theta}^{i}}E_{\boldsymbol{\theta}^{i}},$ (17)
where $i$ is the iteration step, $\eta$ is the learning rate, $\epsilon=10^{-3}$ is taken for numerical stability, and $S$ is the Quantum Fisher information matrix,
$S_{ab}=\frac{\langle\partial_{\theta_{a}}\psi_{\boldsymbol{\theta}}|\partial_{\theta_{b}}\psi_{\boldsymbol{\theta}}\rangle}{\langle\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}-\frac{\langle\partial_{\theta_{a}}\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}{\langle\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}\frac{\langle\psi_{\boldsymbol{\theta}}|\partial_{\theta_{b}}\psi_{\boldsymbol{\theta}}\rangle}{\langle\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}.$ (18)
@@
During initial optimization, the trial wave function resides far from the ground state in parameter space. We generate a small sample of $N=2\times 10^{4}$ points to rapidly estimate $E_{\boldsymbol{\theta}}$ and $\nabla_{\boldsymbol{\theta}}E_{\boldsymbol{\theta}}$ with deliberate coarseness, enabling efficient early-stage convergence. As the DeepQuark wave function approaches convergence, we increase the sample size to $N=4\,\text{--}\,8\times 10^{4}$ to stabilize the optimization process. During the final optimization stage, we save 10 distinct sets of wave function parameters at a 50-step interval. The energy and other physical observables of these wave function configurations are subsequently evaluated using $N\sim 10^{6}$ sample points. The parameter set yielding the lowest energy expectation value is selected as the final ground-state wave function.
Appendix C Results of electron systems
The ground-state energy results of few-electron systems are listed in Table 6. For comparative purposes, we include well-established benchmark results such as the exact solution of $\mathrm{Ps}$, and high-precision variational results for $\mathrm{Ps}^{-}$ using Hylleraas-type wave functions [74] and $\mathrm{Ps}_{2}$ using explicitly correlated Gaussian functions [75, 76].
Table 6: Ground-state energies (in eV) of the few-electron systems. The statistical errors of the DeepQuark results are shown in the parentheses. Results from other methods are shown for comparison. $\mathrm{Ps}$ can be solved exactly while $\mathrm{Ps}^{-}$ and $\mathrm{Ps}_{2}$ are solved using variational method.

### v2

@@ A.3 Applications to other models
The deep neural network (DNN) in the DeepQuark wave function consists of the input features $\boldsymbol{x}=\left(\boldsymbol{r}_{i},\left|\boldsymbol{r}_{i}-\boldsymbol{r}_{j}\right|,\alpha_{c},\alpha_{s},\alpha_{t}\right)$, four hidden fully connected layers and a one-dimensional output $f_{NN}(\boldsymbol{x})$. Here $\boldsymbol{r}_{i},\alpha_{c},\alpha_{s},\alpha_{t}$ are the spatial three-dimensional coordinates in the center of mass frame and the encoded vectors of color, spin and isospin coupled bases, respectively. The number of nodes in each hidden layer and the total number of variational parameters in the DNN for different systems are given in Table SM-VI. We use larger DNNs for larger systems to ensure the flexibility of the wave function.
@@
Table SM-VI: The number of nodes in each hidden layer and the total number of variational parameters in the DNN for different systems.
@@
We optimize the variational parameters $\boldsymbol{\theta}$ by minimizing the energy expectation value. In each iteration step during the optimization process, we use the Metropolis-Hastings Monte Carlo algorithm [84, 83] to generate sample points that are distributed according to the probability
$P(R,\alpha)=\frac{\left|\psi_{\boldsymbol{\theta}}(R,\alpha)\right|^{2}}{\sum_{\alpha}\int dR\left|\psi_{\boldsymbol{\theta}}(R,\alpha)\right|^{2}},$ (S10)
@@
$P=\frac{\left|\psi_{\boldsymbol{\theta}}\left(R^{\prime},\alpha^{\prime}\right)\right|^{2}}{\left|\psi_{\boldsymbol{\theta}}(R,\alpha)\right|^{2}}.$ (S11)
The energy expectation value $E_{\boldsymbol{\theta}}=\frac{\langle\psi_{\boldsymbol{\theta}}|H|\psi_{\boldsymbol{\theta}}\rangle}{\langle\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}$ and its gradient with respect to the parameters $\nabla_{\boldsymbol{\theta}}E_{\boldsymbol{\theta}}$ can be estimated using these sample points.
$E_{\boldsymbol{\theta}}$ $=\frac{\sum_{\alpha\alpha^{\prime}}\int dR\,\psi_{\boldsymbol{\theta}}^{*}(R,\alpha)H_{\alpha\alpha^{\prime}}(R)\psi_{\boldsymbol{\theta}}(R,\alpha^{\prime})}{\sum_{\alpha}\int dR\,|\psi_{\boldsymbol{\theta}}(R,\alpha)|^{2}}$ (S12)
$=\sum_{\alpha\alpha^{\prime}}\int dR\,P(R,\alpha)\psi_{\boldsymbol{\theta}}^{-1}(R,\alpha)H_{\alpha\alpha^{\prime}}(R)\psi_{\boldsymbol{\theta}}(R,\alpha^{\prime})$
@@
$E_{L}(R,\alpha)=\sum_{\alpha^{\prime}}\psi_{\boldsymbol{\theta}}(R,\alpha)^{-1}H_{\alpha\alpha^{\prime}}(R)\psi_{\boldsymbol{\theta}}(R,\alpha^{\prime}).$ (S13)
@@
$\nabla_{\boldsymbol{\theta}}E_{\boldsymbol{\theta}}=2$ $\left[\frac{1}{N}\left(\sum_{n}E_{L}(R_{n},\alpha_{n})\nabla_{\boldsymbol{\theta}}\log\psi_{\boldsymbol{\theta}}(R_{n},\alpha_{n})\right)\right$.- (S14)
@@
$\boldsymbol{\theta}^{i+1}=\boldsymbol{\theta}^{i}-\eta(S+\epsilon I)^{-1}\nabla_{\boldsymbol{\theta}^{i}}E_{\boldsymbol{\theta}^{i}},$ (S15)
where $i$ is the iteration step, $\eta$ is the learning rate, $\epsilon=10^{-2}$ is taken for numerical stability, and $S$ is the Quantum Fisher information matrix,
$S_{ab}=\frac{\langle\partial_{\theta_{a}}\psi_{\boldsymbol{\theta}}|\partial_{\theta_{b}}\psi_{\boldsymbol{\theta}}\rangle}{\langle\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}-\frac{\langle\partial_{\theta_{a}}\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}{\langle\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}\frac{\langle\psi_{\boldsymbol{\theta}}|\partial_{\theta_{b}}\psi_{\boldsymbol{\theta}}\rangle}{\langle\psi_{\boldsymbol{\theta}}|\psi_{\boldsymbol{\theta}}\rangle}.$ (S16)
@@
During initial optimization, the trial wave function resides far from the ground state in parameter space. We generate a small sample of $N=2\times 10^{4}$ points to rapidly estimate $E_{\boldsymbol{\theta}}$ and $\nabla_{\boldsymbol{\theta}}E_{\boldsymbol{\theta}}$ with deliberate coarseness, enabling efficient early-stage convergence. As the DeepQuark wave function approaches convergence, we increase the sample size to stabilize the optimization process. During the final optimization stage, we save 10 distinct sets of wave function parameters at a 50-step interval. The energy and other physical observables of these wave function configurations are subsequently evaluated using $N\sim 10^{6}$ sample points. The parameter set yielding the lowest energy expectation value is selected as the final ground-state wave function.
For each multiquark system, it only take merely $\mathcal{O}(1)$ A100 GPU-hours to get ground state solution. Although direct comparison between phenomenological models and ab initio simulations is improper, we provide lattice QCD ensemble generation costs solely to contextualize computational scales. For example, it was reported the CLQCD collaboration spent over $6\times 10^{5}$ NVIDIA A100 GPU-hours to generate only part of the planned ensembles [86].
Supplemental Material C Results of electron systems
The ground-state energy results of few-electron systems are listed in Table SM-VII. For comparative purposes, we include well-established benchmark results such as the exact solution of $\mathrm{Ps}$, and high-precision variational results for $\mathrm{Ps}^{-}$ using Hylleraas-type wave functions [82] and $\mathrm{Ps}_{2}$ using explicitly correlated Gaussian functions [79, 87].
Table SM-VII: Ground-state energies (in eV) of the few-electron systems. The statistical errors of the DeepQuark results are shown in the parentheses. Results from other methods are shown for comparison. $\mathrm{Ps}$ can be solved exactly while $\mathrm{Ps}^{-}$ and $\mathrm{Ps}_{2}$ are solved using variational method.